# Arxiv preprint

### v1

@@ Abstract
Generative AI can turn scientific articles into narratives for diverse audiences, but evaluating these stories remains challenging. Storytelling demands abstraction, simplification, and pedagogical creativity—qualities that are not often well-captured by standard summarization metrics. Meanwhile, factual hallucinations are critical in scientific contexts, yet, detectors often misclassify legitimate narrative reformulations or prove unstable when creativity is involved. In this work, we propose StoryScore, a composite metric for evaluating AI-generated scientific stories. StoryScore integrates semantic alignment, lexical grounding, narrative control, structural fidelity, redundancy avoidance, and entity-level hallucination detection into a unified framework. Our analysis also reveals why many hallucination detection methods fail to distinguish pedagogical creativity from factual errors, highlighting a key limitation: while automatic metrics can effectively assess semantic similarity with original content, they struggle to evaluate how it is narrated and controlled.
@@ 1 Introduction
Generative AI in general and Large language models (LLMs) in particular can summarize and rephrase complex content, presenting it in a narrative form that makes it accessible to non-experts [teleki-etal-2025-survey]. Despite these advances, LLMs frequently struggle to appropriately adapt tone, level of detail, and stylistic choices to diverse audiences and communication objectives, regardless of the amount of contextual instructions and roles than one can use when prompting. In addition, LLMs are prone to hallucinations, eloquently stating unsupported claims, which pose a significant risk in scientific communication [ji2023survey].
Recent works are exploring scientific storytelling as a way to generate structured and engaging narratives that explain scientific content and papers while tailoring explanations to specific personas [roschelle2025generative, LI2023101311]. However, evaluating such narratives presents challenges that differ fundamentally from those associated with traditional summarization tasks. Persona-oriented storytelling deliberately introduces abstraction, metaphor, and contextualization to enhance accessibility. These transformations are desirable for communication but significantly complicate evaluation. In particular, assessing factual grounding becomes non-trivial when the generated text is not expected to closely mirror the source document.
While the development of scientific storytelling systems is gaining momentum, the question of how to evaluate the resulting narratives remains largely open. In particular:
@@
In this work, we address these questions in the context of designing a system that transform scientific papers into audience-adapted stories. We make two main contributions:
@@
The remainder of the paper is organised as follows. After reviewing the relevant literature in Section 2, we introduce the background use case in Section 3. We detail the StoryScore metric in Section 4. Next, we discuss the limitations of existing hallucination detection methods for our use case in Section 5. We present the findings of our experiments in Section 6. Finally, we conclude and outline some future work in Section 7. The code [^1] https://anonymous.4open.science/r/scientist-storyteller-2026 and data [^2] https://zenodo.org/records/18301594 used in this work are publicly available.
@@ 2 Related Work
Evaluating the quality of AI-generated stories or summaries has also advanced beyond simple word-overlap scores. Traditional metrics like ROUGE [lin-2004-rouge] measure n-gram overlap with reference summaries, but they correlate poorly with human judgments [afzal2023challenges, scialom-etal-2021-questeval, chhun-etal-2022-human]. Semantic metrics such as BERTScore [zhang2019bertscore] and MoverScore [zhao-etal-2019-moverscore] use contextual embeddings to better capture semantic alignment and compare meaning rather than surface words. They achieve high correlation with human ratings on summarization and other tasks, but remain largely insensitive to structural degradation, redundancy, and discourse-level control failures, which are critical in storytelling. This gap is amplified in multi-section stories, where discourse-level issues (e.g., global coherence, redundancy, or instruction leakage) strongly affect perceived quality while remaining weakly captured by similarity-based metrics. Other methods, such as QuestEval [scialom-etal-2021-questeval] and QAGS [wang-etal-2020-asking] ask questions about a summary and its source to detect factual inconsistencies.
@@
Hallucination detection has been studied extensively in summarization and question answering, using entity-based checks, retrieval and entailment methods, and LLM-based judges [huang2025survey, ji2023survey, li2025generation]. Systems can use RAG-based techniques on specific documents or on the web, like the widely adopted GPTZero [^3] https://gptzero.me/. Notably, recent empirical work shows pitfalls in automated detection. For example, they show that ROUGE and similar metrics can be easily fooled – e.g. repetitively duplicating correct content (adding length) artificially inflates ROUGE scores, even though no new facts are added [janiak-etal-2025-illusion]. In addition, common approaches often assume strict fidelity to the source. In persona-adaptive storytelling, legitimate reformulations and contextual expansions are expected. Thus, detectors can over-penalize creativity or behave unstably. We will critically analyse existing methods in Section 5.
3 Use Case: An AI-Driven Pipeline for Scientific Story Generation
The metrics proposed in this paper are grounded in a concrete use case: generating persona-adapted explanatory stories from scientific papers. We employ a two-stage pipeline in which a Splitter module produces a five-section outline from a scientific paper, and a Storyteller module expands each section into a persona-conditioned narrative. In our implementation, the Splitter uses the Qwen2.5–7B [qwen2025qwen25technicalreport] model for stable structured output, while the Storyteller relies on the Qwen2.5–32B model for long-context generation, persona adherence, and reduced hallucinations. Specific prompts constrain the output structure and discourage the invention of entities, numeric values, or references, while allowing simplification and contextualization.
We conducted experiments on a curated dataset of 62 scientific papers paired with 190 human-written explanatory stories across three domains (Entertainment, Accessibility, Health) and eight personas. From this corpus, we derive 496 outline instances for training the Splitter and 818 story-level instances (4,090 section texts) for training the Storyteller, using standard train/validation/test splits. The test set is used to generate persona-adapted stories that serve as evaluation material for StoryScore and for the comparative analysis of hallucination detection methods detailed below.
@@ 4 The StoryScore metric for Generated Story Evaluation
For this reason, we propose StoryScore, a composite metric aggregating semantic similarity, lexical grounding, structural fidelity, fluency, and hallucination control into a single score in the range $[0,1]$. Following extensive experimentation on a real use case, including a multi-stage analysis of hallucination detection explained later, it is defined as:
@@
This formulation captures the most stable and reliable indicators, maintaining a balanced representation of narrative quality dimensions. The weights were selected heuristically after exploratory adjustments guided by manual qualitative inspection, with the goal of producing an aggregate score consistent with human qualitative judgments. The components of StoryScore are summarised in Table 1.
@@
StoryScore Composite Global weighted quality score
Context recall or Article coverage quantifies the amount of content from the original article that is reflected in the generated story, measured as the proportion of word-level lexical tokens from the article that also appear in the narrative (article-centered coverage). Tokens are defined as lowercased words, excluding punctuation and stopwords. Unlike metrics that rely solely on lexical overlap, contextual recall serves as a proxy for content coverage, quantifying the extent to which the generated text captures the key concepts, terminology, and salient tokens from the original article. Formally, it is defined as:
@@
where $T_{\mathrm{paper}}$ and $T_{\mathrm{story}}$ denote the sets of tokens extracted from the paper and the generated story. Higher values indicate stronger grounding in the source material, but also a vocabulary that is more faithful to the vocabulary used in the paper and less persona-adapted for an audience with low expertise. We intentionally adopt a simple set-based formulation to obtain a transparent and stable signal, acknowledging that this choice favours lexical grounding over semantic abstraction and may penalize aggressive simplification.
BERTScore [zhang2019bertscore] measures semantic faithfulness by comparing contextual embeddings of the generated story with those of the source article. BERTScore aligns tokens from the hypothesis to the most semantically similar tokens in the reference using cosine similarity in the embedding space. For each token, the F1-score are computed by aggregating the highest-scoring matches. Formally, let $H$ and $R$ be the embedding sets for the hypothesis (story) and reference (paper), BERTScore is defined as:
@@
Prompt Cleanliness measures the extent to which the generated story is free of prompt-related artifacts and instruction leaks, such as residual system directives, role indicators, formatting constraints, or meta-level indications. These artifacts indicate a failure of narrative abstraction, where the model collapses from storytelling to instruction-following behaviour, severely degrading readability. The generated text is analysed line by line and sentence by sentence using a set of high-precision regular expression patterns anchored to sentence or line boundaries. A contamination score $C$ is computed as:
@@
  • $N_{\mathrm{line}}$ is the number of lines beginning with explicit chat-role or instruction markers (e.g., Human:, Assistant:, Rules:);
@@
$\mathit{PromptCleanliness}=1-\min(1,C).$ (5)
Title Coverage evaluates whether the generated story preserves the section structure of the original outline. Let $\mathcal{O}=\{o_{1},\dots,o_{5}\}$ denote the target section titles and $\mathcal{G}=\{g_{1},\dots,g_{5}\}$ the generated ones. We compute a soft similarity score and average it across sections:
$\mathit{TitleCoverage}=\frac{1}{n}\sum_{i=1}^{n}\mathrm{sim}\bigl(\mathrm{norm}(g_{i}),\,\mathrm{norm}(o_{i})\bigr),$ (6)
@@
$\mathit{RedundancyRate}=\frac{\max_{g\in\mathcal{G}_{n}}\mathrm{freq}(g)}{|\mathcal{G}_{n}|},\qquad\mathit{NoRedundancy}=1-\mathit{RedundancyRate}.$ (7)
@@
No Hallucination quantifies the inclusion of entities that are not supported by the source material. After a post-processing normalization step applied to extracted entities (e.g. lowercasing, removal of possessive suffixes), we perform a NER-based comparison: the generated story is analyzed using SpaCy [spacy2020] to extract entities of type PERSON and ORG, which are then matched against the entities detected in the source paper. The restriction to these entity types is deliberate: in our manual qualitative inspection of generated outputs, hallucinations most frequently manifested as invented author names and institutional affiliations, making PERSON/ORG the most relevant entity types for a stable automatic signal. Any entity appearing in the story but absent from the source material is treated as a hallucination. A broader discussion of alternative hallucination detectors, along with their limitations and the decision of why this is the approach used, is provided in Section 5. Formally, let $\mathit{GeneratedEntities}$ be the set of PERSON/ORG entities extracted from the story, and $\mathit{HallucinatedEntities}$ the subset of those not found in the paper, the score is defined as:
$\mathit{NoHallucination}=1-\frac{\lvert\mathit{HallucinatedEntities}\rvert}{\lvert\mathit{GeneratedEntities}\rvert}$. (8)
A higher value indicates stronger factual alignment with the source document.
@@ 5 Hallucination Detection in Scientific Storytelling
A key challenge in AI-based story generation from source documents is assessing whether generated narratives remain faithful to the underlying material without introducing unsupported entities or claims. Unlike summarization, storytelling intentionally involves simplification, contextualization, and narrative adaptation, producing an expected and desirable creative divergence. The core challenge is in distinguishing legitimate narrative abstraction from true factual hallucination, a distinction that existing hallucination detection methods (largely designed for summarization) are not well equipped to make.
@@
In this approach, we leverage SpaCy to only detect named entities of types PERSON and ORG, as described in Section 4 and Equation 8. This significantly reduces the number of irrelevant candidates and detects genuinely invented organisations or people introduced by the model. Nevertheless, NER remained a shallow signal: it can detect incorrect affiliations, but fails to identify deeper factual inconsistencies, e.g. wrong scientific claims, invented datasets, or unsupported causal statements. We specifically focus on PERSON/ORG because our qualitative analysis repeatedly found fabricated authors and affiliations to be the most frequent and disruptive hallucination pattern in our stories.
@@
MIRAGE is an open-source library for hallucination detection based on rewriting the same passage multiple times and measuring the stability of the appearing concepts [10.1145/3746252.3761644]. If an idea disappears or mutates across rewrites, MIRAGE treats it as hallucinated. Although effective in summarization, this approach does not transfer well to persona-oriented storytelling. As revealed by our experiments, MIRAGE consistently and incorrectly flags explanatory metaphors (e.g., introducing a system by analogy with a familiar real-world process that is not explicitly described in the paper) and rephrasing for non-expert audiences (e.g., replacing technical terminology with higher-level conceptual descriptions). This indicates a misalignment between MIRAGE’s operational definition of fidelity, centred on literal grounding, and the goals of narrative systems that intentionally simplify the source material.
@@
A first experiment conducted with the Qwen–7B model [qwen2025qwen25technicalreport] has shown two systematic patterns:
@@
    Missed hallucinations: Some fabricated facts (e.g. invented institutional affiliations) are left undetected.
@@
In short, the judge sometimes hallucinates hallucinations, making it too unreliable to serve as a scoring component.
@@
This approach combines symbolic robustness (entity extraction), semantic flexibility (retrieval-based context), and local grounding (sentence-level comparison), but is difficult to calibrate: if the threshold is too low, subtle hallucinations pass undetected; if too high, creative paraphrases are over-penalised.
We qualitatively analysed the Hybrid Hallucination Detector (HHD) across three papers and their generated story sections. Manual inspection showed that false positives dominate the detector’s output, mainly due to pedagogical reformulations that are conceptually faithful but not literally supported by retrieved sentences. For example, the following excerpt is from a generated story about [gao2024training]:
@@
In this case, HERMES is the name of a framework, introduced through a creative but semantically correct analogy, yet incorrectly flagged as hallucinated. Conversely, false negatives occurred when the retrieval step returned semantically adjacent but non-supportive contexts. Taking another example from the same story:
@@
The original paper used the acronym FM to refer to Foundation Models, but not Flash Memory, resulting in an undetected hallucination.
Overall, the method proved too sensitive to surface mismatch and too permissive to semantic drift. Because these two error types move in opposite directions, tuning the threshold did not lead to a stable operating point, making HHD unsuitable as a scoring component for StoryScore.
@@
Table 2: Comparison of explored hallucination detection methods.
Table 2 summarizes our findings on different hallucination detection techniques. From our experiments, employing these techniques on the use case, two key outcomes emerged. First, The most operationally stable detector was the simplest one, namely the NER-Based detection, combined with regex normalisation. In particular, it was the only method that: 1. remained stable across papers, showing consistent behaviour across different documents, 2. did not penalise metaphors, 3. detected genuine fabricated entities, and 4. integrated cleanly with a software pipeline, without the need of external and costly API calls. Second, hallucination must have a reduced weight in the overall metric. Since all methods exhibited weaknesses, either too naive (NER), too strict (MIRAGE), or too unstable (LLM judge), its contribution to StoryScore was reduced from 0.2 (which was initially planned) to 0.1. This ensures that hallucination detection informs the metric without overshadowing the other reliable qualities.

### v2

# R. Campos, A. Jorge, A. Jatowt, S. Bhatia, M. Litvak (eds.): Proceedings of the Text2Story’26 Workshop, Delft (The Netherlands), 29-March-2026

@@ Abstract
Generative AI can turn scientific articles into narratives for diverse audiences, but evaluating these stories remains challenging. Storytelling demands abstraction, simplification, and pedagogical creativity – qualities that are not often well-captured by standard summarization metrics. Factual hallucinations are critical in scientific contexts, yet, detectors often misclassify legitimate narrative reformulations. In this work, we propose StoryScore, a composite metric for evaluating AI-generated scientific stories. StoryScore integrates semantic alignment, lexical grounding, narrative control, structural fidelity, redundancy avoidance, and entity-level hallucination detection into a unified framework. Our analysis also reveals why many hallucination detection methods fail to distinguish pedagogical creativity from factual errors, highlighting a key limitation: while automatic metrics can effectively assess semantic similarity with original content, they struggle to evaluate how it is narrated and controlled.
@@ 1 Introduction
Generative AI and Large language models (LLMs) can summarize and rephrase complex content, presenting it in a narrative form that makes it accessible to non-experts [teleki-etal-2025-survey]. Despite these advances, LLMs frequently struggle to appropriately adapt tone, level of detail, and stylistic choices to diverse audiences and communication objectives. In addition, LLMs are prone to hallucinations, eloquently stating unsupported claims, which pose a significant risk in scientific communication [ji2023survey].
Recent works are exploring scientific storytelling as a way to generate structured and engaging explanation of scientific content and papers, tailored to specific personas [roschelle2025generative, LI2023101311, sillano2026personas]. Evaluating such narratives presents challenges fundamentally different from those faced in traditional summarisation tasks. Storytelling deliberately introduces abstraction, metaphor, and contextualization to enhance accessibility. At the same time, this complicates evaluation: assessing factual grounding becomes non-trivial when the generated text is not expected to closely mirror the source document. So, the question of how to evaluate generated narratives remains largely open. In particular:
@@
In this work, we address these questions in the context of designing a system that transforms scientific papers into audience-adapted stories. We make two main contributions:
@@
The metrics are grounded in a concrete use case: a two-stage pipeline (a Splitter defining the outline and a Storyteller in charge of the writing) that transforms scientific papers into persona-adapted stories.
The code used in this work is available at https://github.com/AlexArgese/ai-scientist-storyteller.
@@ 2 Related Work
Evaluating the quality of AI-generated summaries has also advanced beyond simple word-overlap scores. Traditional metrics like ROUGE [lin-2004-rouge] measure n-gram overlap with reference summaries, but they correlate poorly with human judgments [afzal2023challenges, scialom-etal-2021-questeval, chhun-etal-2022-human]. Semantic metrics such as BERTScore [zhang2019bertscore] and MoverScore [zhao-etal-2019-moverscore] use contextual embeddings to better capture semantic alignment and compare meaning rather than surface words. They achieve high correlation with human ratings on summarization, but remain largely insensitive to structural degradation, redundancy, and discourse-level control failures. This gap is amplified in multi-section stories, where discourse-level issues (e.g., global coherence or instruction leakage) strongly affect perceived quality while remaining weakly captured by similarity-based metrics. Other methods, such as QuestEval [scialom-etal-2021-questeval] and QAGS [wang-etal-2020-asking], ask questions about a summary and its source to detect inconsistencies.
@@
Hallucination detection has been studied extensively in summarization and question answering, using entity-based checks, retrieval and entailment methods, and LLM-based judges [ji2023survey, huang2025survey, li2025generation]. Systems can use RAG-based techniques on specific documents or on the web, like the widely adopted GPTZero [^1] https://gptzero.me/. Recent empirical work shows pitfalls in automated detection. For example, they show that ROUGE and similar metrics can be easily fooled – e.g. repetitively duplicating correct content (adding length) artificially inflates ROUGE scores, even though no new facts are added [janiak-etal-2025-illusion]. In addition, common approaches often assume strict fidelity to the source. In persona-adaptive storytelling, legitimate reformulations and contextual expansions are expected. Thus, detectors can over-penalize creativity or behave unstably. We will critically analyse existing methods in Section 4.
@@ 3 The StoryScore metric for Generated Story Evaluation
For this reason, we propose StoryScore, a composite metric aggregating semantic similarity, lexical grounding, structural fidelity, fluency, and hallucination control into a single score in the range $[0,1]$. Following extensive experimentation on a real use case, it is defined as:
@@
This formulation captures the most stable and reliable indicators, maintaining a balanced representation of narrative quality dimensions. The weights were selected heuristically after exploratory adjustments guided by manual qualitative inspection, with the goal of producing an aggregate score. The components of StoryScore are summarised in Table 1.
@@
Context recall or Article coverage quantifies the amount of content from the original article that is reflected in the generated story, measured as the proportion of word-level lexical tokens from the article that also appear in the narrative (article-centered coverage). Tokens are defined as lowercased words, excluding punctuation and stopwords. Unlike metrics that rely solely on lexical overlap, contextual recall serves as a proxy for content coverage. Formally, it is defined as:
@@
where $T_{\mathrm{paper}}$ and $T_{\mathrm{story}}$ denote the sets of tokens extracted from the paper and the generated story. Higher values suggest a stronger connection to the original source material but also a vocabulary more faithful to the paper’s language and less suited for a less knowledgeable audience. We deliberately adopt a simple set-based formulation to achieve a transparent and stable signal. This choice prioritises lexical grounding over semantic abstraction and may therefore disadvantage aggressive simplification.
BERTScore [zhang2019bertscore] measures semantic faithfulness by comparing contextual embeddings of the generated story with those of the source article. BERTScore aligns tokens from the hypothesis to the most semantically similar tokens in the reference using cosine similarity in the embedding space. For each token, the F1-score is computed by aggregating the highest-scoring matches. Formally, let $H$ and $R$ be the embedding sets for the hypothesis (story) and reference (paper), BERTScore is defined as:
@@
Prompt Cleanliness measures the extent to which the generated story is free of prompt-related artifacts and instruction leaks, such as residual system directives, role indicators, formatting constraints, or meta-level indications. These artifacts indicate a failure of narrative abstraction, where the model collapses from storytelling to instruction-following behaviour, severely degrading readability. The generated text is analysed line by line and sentence by sentence using a set of regular expression patterns that we empirically found as relevant. A contamination score $C$ is computed as:
@@
  • $N_{\mathrm{line}}$ is the number of lines beginning with explicit instruction markers (e.g., Human:, Rules:);
@@
$\mathit{PromptCleanliness}=1-\min(1,C)$ (5)
Title Coverage evaluates whether the generated story preserves the section structure of the target outline [^2] In our use case, this is the outline generated by the Splitter module, as in Section 1.. Let $\mathcal{O}=\{o_{1},\dots,o_{5}\}$ denote the target section titles and $\mathcal{G}=\{g_{1},\dots,g_{5}\}$ the generated ones. We compute a soft similarity score and average it across sections:
$\mathit{TitleCoverage}=\frac{1}{n}\sum_{i=1}^{n}\mathrm{sim}\bigl(\mathrm{norm}(g_{i}),\,\mathrm{norm}(o_{i})\bigr)$ (6)
@@
$\mathit{RedundancyRate}=\frac{\max_{g\in\mathcal{G}_{n}}\mathrm{freq}(g)}{|\mathcal{G}_{n}|},\qquad\mathit{NoRedundancy}=1-\mathit{RedundancyRate}$ (7)
@@
No Hallucination quantifies the inclusion of entities that are not supported by the source material. After a post-processing normalization step applied to extracted entities (e.g. lowercasing, removal of possessive suffixes), we perform a NER-based comparison: the generated story is analyzed using SpaCy [spacy2020] to extract entities of type PERSON and ORG, which are then matched against the entities detected in the source paper. The restriction to these entity types is deliberate: in our manual qualitative inspection of generated outputs, hallucinations most frequently manifested as invented author names and institutional affiliations, making PERSON/ORG the most relevant entity types for a stable automatic signal. Any entity appearing in the story but absent from the source material is treated as a hallucination. A broader discussion of alternative hallucination detectors is provided in Section 4. Formally, let $\mathit{GeneratedEntities}$ be the set of PERSON/ORG entities extracted from the story, and $\mathit{HallucinatedEntities}$ the subset of those not found in the paper, the score is defined as:
$\mathit{NoHallucination}=1-\frac{\lvert\mathit{HallucinatedEntities}\rvert}{\lvert\mathit{GeneratedEntities}\rvert}$ (8)
@@ 4 Hallucination Detection in Scientific Storytelling
Unlike summarization, storytelling intentionally involves simplification, contextualization, and narrative adaptation, producing an expected and desirable creative divergence. The core challenge is in distinguishing legitimate narrative abstraction from true factual hallucination, a distinction that existing hallucination detection methods (largely designed for summarization) are not well equipped to make.
The following analysis is grounded in a qualitative inspection of AI-generated stories produced by the pipeline on a representative subset of papers from the corpus. Each hallucination detection method was applied to the same generated stories and its outputs were manually examined to identify systematic failure modes specific to persona-adaptive scientific storytelling.
@@
We leverage SpaCy to only detect named entities of types PERSON and ORG, as described in Section 3. This significantly reduces the number of candidates and detects genuinely invented organisations or people introduced by the model. Nevertheless, NER remained a shallow signal: it can detect incorrect affiliations, but fails to identify deeper factual inconsistencies, e.g. wrong scientific claims, invented datasets, or unsupported causal statements. We specifically focus on PERSON/ORG because our qualitative analysis repeatedly found fabricated authors and affiliations to be the most frequent and disruptive hallucination pattern in our stories.
@@
MIRAGE is a library for hallucination detection based on rewriting the same passage multiple times and measuring the stability of the appearing concepts [10.1145/3746252.3761644]. If an idea disappears or mutates across rewrites, MIRAGE treats it as hallucinated. Although effective in summarization, this approach does not transfer well to persona-oriented storytelling. In our experiments, MIRAGE consistently and incorrectly flags explanatory metaphors (e.g., introducing a system by analogy with a familiar real-world process that is not explicitly described in the paper) and rephrasing for non-expert audiences (e.g., replacing technical terminology with higher-level conceptual descriptions). This indicates a misalignment between MIRAGE’s operational definition of fidelity, centred on literal grounding, and the goals of narrative systems that intentionally simplify the source material.
@@
A first experiment conducted with the Qwen2.5-7B model [qwen2025qwen25technicalreport] has shown two systematic patterns:
@@
    Missed hallucinations: Some fabricated facts (e.g. invented affiliations) are left undetected.
@@
In short, the judge sometimes hallucinates hallucinations, making it too unreliable.
@@
This approach combines symbolic robustness (entity extraction), semantic flexibility (retrieval-based context), and local grounding (sentence-level comparison), but is difficult to calibrate.
Manual inspection on generated stories showed that false positives dominate the detector’s output, mainly due to pedagogical reformulations that are conceptually faithful but not literally supported by retrieved sentences. For example, the following excerpt is from a generated story about [gao2024training]:
@@
In this case, HERMES is the name of a framework, introduced through a creative but semantically correct analogy, yet incorrectly flagged as hallucinated. Conversely, false negatives occurred when the retrieval step returned semantically adjacent but non-supportive contexts. Another example from the same story fails to correctly assign the acronym FM to Foundation Models, resulting in the following undetected hallucination:
@@
Because these two error types move in opposite directions, tuning the threshold did not lead to a stable operating point, making HHD unsuitable as a component for StoryScore.
Table 2: Comparison of explored hallucination detection methods.
@@
Table 2 summarizes our findings on different hallucination detection techniques, that lead to two key outcomes. First, The most operationally stable detector was the simplest one, namely the NER-Based detection, combined with regex normalisation. In particular, it was the only method that: 1. remained stable across papers, showing consistent behaviour across different documents, 2. did not penalise metaphors, 3. detected genuine fabricated entities, and 4. integrated cleanly with a software pipeline, without the need of external and costly API calls. Second, being either too naive (NER), too strict (MIRAGE), or too unstable (LLM judge), the hallucination metric must have a reduced weight in the overall metric (limited to 10%) to inform without overshadowing the other reliable qualities.